### v1

# Shedding Light on Gravity: Black Hole Shadows and Lensing Signatures in Lorentz Gauge Theory

Abstract
Recent advances, including gravitational wave detections and imaging of black hole shadows, have strongly validated general relativity. Nevertheless, ongoing cosmological observations suggest potential limitations of general relativity, spurring interest in modified theories of gravity. This study explores Lorentz gauge theory, an alternative gravitational framework offering promising solutions to longstanding conceptual issues in quantum gravity and cosmology. By analyzing black hole shadow structures and gravitational lensing effects—both weak and strong deflection regimes—we highlight unique observational signatures of Lorentz gauge gravity. Our findings provide valuable tools for future observational tests, potentially distinguishing these modified gravity models from general relativity and advancing our understanding of spacetime geometry and fundamental gravitational interactions.
Black holes; Weak deflection angle; Black hole shadow; Strong lensing; Lorentz gauge theory.
pacs:
95.30.Sf, 04.70.-s, 97.60.Lf, 04.50.+h
I Introduction
General relativity (GR), Einstein’s revolutionary theory, interprets gravity as a manifestation of curved spacetime, dynamically described by a metric tensor. Among its most extraordinary predictions is the existence of black holes—regions in spacetime where gravity is so intense that nothing, not even light, can escape Wald (1984); Misner et al. (1973); Virbhadra and Ellis (2000a, 2002); Claudel et al. (2001); Adler and Virbhadra (2022). The validity and precision of GR have been spectacularly confirmed through recent landmark observations, including gravitational wave detections by LIGO and Virgo collaborations et al. (2016, 2019a) and high-resolution imaging of black hole shadows by the Event Horizon Telescope (EHT) collaboration et al. (2019b, 2022).
Nevertheless, current cosmological observations revealing accelerated cosmic expansion and notable tensions in cosmological parameters have motivated significant interest in exploring modifications to the standard metric dynamics et al. (2012a). Such theoretical developments encompass alternative geometric frameworks and formulations of gravity, moving beyond the traditional metric-based description Sotiriou and Faraoni (2010). A critical conceptual challenge for classical GR is its inherent reliance on a non-degenerate metric; notably, the theory lacks a natural zero-metric ”ground state,” a deficiency linked to significant theoretical difficulties, particularly its non-renormalizability in quantum formulations Weinberg (1979).
In addressing these fundamental issues, alternative formulations of gravity, such as Plebański’s chiral description, provide compelling insights Plebański (1977). Plebański’s framework expresses Einstein’s equations solely through chiral components of the Lorentz group, employing a triple of two-forms and a chiral connection, thus offering a metric-independent initial description of gravity. By imposing suitable conditions, this approach reconstructs spacetime metrics dynamically, underpinning prominent quantum gravity proposals Rovelli (2004); Ashtekar and Lewandowski (2004).
Expanding upon these conceptual advancements, the Lorentz gauge theory (LGT) emerges by gauging the Lorentz invariance of special relativistic parameterized fields, identifying a canonical clock field—the ”khronon”—as fundamental in defining spacetime geometry et al. (2017, 2006). This innovative approach inherently resolves several longstanding conceptual problems of time in quantum gravity, embedding the metric construction within the chiral Lorentz connection itself. The resulting theory elegantly integrates matter fields, naturally including spinor fields and Yang-Mills interactions, without extraneous constraints et al. (2013). Recent Hamiltonian analyses have identified that only the purely chiral curvature action extends effectively to GR, whereas the action incorporating total curvature represents a topological theory devoid of local degrees of freedom Smolin (2004). This intriguing result points towards promising pathways for quantum gravity formulations.
Given these compelling theoretical motivations, further detailed studies of classical solutions are vital. Gravitational lensing phenomena, particularly weak and strong deflection angles, provide powerful observational tests for these gravity theories Weinberg (1972); Virbhadra and Ellis (2000b). The weak deflection angle regime, typically involving minor bending of light around massive objects, serves as an important testbed for GR and its alternatives Keeton and Petters (2005). Conversely, the strong deflection angle regime, encountered in proximity to compact massive objects like black holes, amplifies gravitational lensing effects, offering critical opportunities to detect subtle deviations from standard GR predictions Bozza (2002a); et al. (2012b).
Additionally, the observational characterization of black hole shadows, as impressively demonstrated by the EHT’s imaging of supermassive black holes M87* and Sgr A*, has opened a novel observational window into spacetime structure and gravity theories et al. (2019b, 2022). These shadows result from photon spheres—regions where gravitational fields compel photons into unstable circular orbits, producing a distinctive dark silhouette against surrounding luminous emission Cunha and Herdeiro (2018); Cardoso and Pani (2019). Precision measurements of shadow size and shape thus serve as sensitive probes of gravity theories, potentially revealing deviations indicative of new physics Bambi and Freese (2009); Johannsen and Psaltis (2010a). Observations from the Event Horizon Telescope (EHT), notably imaging the shadow of supermassive black holes M87* and Sgr A*, have stimulated considerable research interest in testing gravity theories, investigating black holes’ scalar hair, and probing the no-hair theorem Vagnozzi and Visinelli (2019); Cardoso and Gualtieri (2016); Khodadi et al. (2020, 2021); Afrin et al. (2021); Johannsen and Psaltis (2010b). Additionally, recent analytical studies have extensively examined how environmental factors like plasma distribution influence black hole shadows and gravitational lensing, yielding critical implications for astrophysical observations Perlick et al. (2015); Bisnovatyi-Kogan and Tsupko (2010, 2018); Perlick and Tsupko (2022).
Modified gravity theories, higher-dimensional models, and nonlinear electrodynamics have also been extensively explored through black hole shadow observations, constraining theoretical parameters effectively Kuang et al. (2022); Belhaj et al. (2021); Pantig et al. (2022); Puliçe et al. (2023); Allahyari et al. (2020); Okyay and Övgün (2022); Övgün (2021). Investigations into exotic objects such as naked singularities and rotating traversable wormholes further highlight the diverse phenomenological consequences visible in gravitational lensing and shadow imaging Shaikh et al. (2019); Joshi et al. (2020); Gyulchev et al. (2018).
Additionally, studies involving shadows of rotating non-Kerr black holes, Einstein-Gauss-Bonnet gravity models, and various scalar-tensor theories continue to refine our understanding of black hole spacetime geometry and potential deviations from General Relativity Atamurotov et al. (2013); Zeng et al. (2020); Cunha and Herdeiro (2020); Cunha et al. (2017); Wei et al. (2019); Jafarzade et al. (2021); Anacleto et al. (2021). Furthermore, recent insights into quasinormal modes and greybody factors provide additional tools for distinguishing theoretical predictions from observational data, crucial for future gravitational wave and electromagnetic observational campaigns Gogoi et al. (2024); Araújo Filho (2025); Wei et al. (2015); Konoplya (2019).
Motivated by these theoretical insights and observational advances, this paper focuses on examining black hole solutions, their shadow structures, and gravitational deflection angles within the LGT framework. Such analyses may illuminate distinctive signatures that differentiate modified theories of gravity from classical GR, thereby contributing significantly to the quest for a deeper understanding of gravitational phenomena.
II Black holes in LGT
In this framework, the dynamical fields include a scalar field $\phi^{\alpha}$, which transforms under the fundamental representation of the Lorentz group, and a connection 1‐form $\omega^{a}{}_{b}$ in the adjoint representation. The connection naturally introduces the covariant exterior derivative $D$, and its associated curvature 2‐form is defined as $R^{a}{}_{b}=d\omega^{a}{}_{b}+\omega^{a}{}_{c}\wedge\omega^{c}{}_{b}.$ This expression encapsulates the nontrivial geometry of the SO(1,3) principal bundle over the 4-dimensional manifold $M$ Koivisto (2023); Koivisto and Zheng (2025); Wiesendanger (2019). In essence, the curvature $R^{a}{}_{b}$ measures the failure of the connection to be locally pure gauge, and it plays a central role in the gravitational dynamics of the theory.
For vacuum solutions, the field equations simplify considerably, leading to explicit forms for the metric functions. In particular, the lapse function $f(r)$, which can be interpreted as the gravitational potential, is given by Koivisto and Zheng (2025)
$f(r)=\frac{1}{A^{2}_{0}}-\frac{m_{S}}{4\pi m^{2}_{P}r}\,,$ (1)
where $m_{P}$ and $m_{S}$ are mass parameters, and $A_{0}$ is the connection coefficient. This equation not only determines the redshift of clocks (through the time component of the metric) but also implicitly affects the spatial geometry by setting a scale for the curvature. Notice that the first term, $1/A_{0}^{2}$, represents a constant offset that directly modifies the gravitational potential, while the second term shows the familiar $1/r$ falloff modulated by the mass parameters.
Substituting this lapse function into the general static, spherically symmetric metric ansatz yields
$ds^{2}=-\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)dt^{2}+\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)^{-1}dr^{2}$
$+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (2)
as demonstrated in Koivisto and Zheng (2025). This metric clearly illustrates how the parameter $A_{0}$ not only scales the time component but also rescales the radial part of the geometry. In effect, the connection coefficient $A_{0}$ plays a dual role, modulating the gravitational redshift and the spatial curvature simultaneously.
A particularly illuminating case arises when one sets $A_{0}=\pm 1$. In this situation, the metric reduces to the familiar Schwarzschild solution, thereby recovering the standard event horizon at $r=2M$ (after redefining the combination $m_{S}/(4\pi m_{P}^{2})$ as $2M$, with $M$ representing the mass of the black hole). This equivalence underscores that $A_{0}$ acts as a scaling parameter: deviations from $|A_{0}|=1$ yield modified spacetimes. Specifically, for $A_{0}>1$, the effective radius of the event horizon increases, and the lapse function $f(r)$ decreases, suggesting that the black hole appears “inflated” compared to its Schwarzschild counterpart. Conversely, if $A_{0}<1$, the event horizon contracts, potentially leading to radical changes in the black hole’s structure and, in extreme cases, to the complete disappearance of the horizon.
The equations not only define the geometric structure of the theory but also highlight the significant role played by the connection parameter $A_{0}$ in determining the physical properties of static, spherically symmetric spacetimes. The interplay between the gravitational potential (via $f(r)$) and the scaling parameter $A_{0}$ provides deeper insights into how modifications in the connection can lead to deviations from classical solutions, such as the Schwarzschild metric.
The work in Abdelqader and Lake (2015) introduces a set of curvature scalars designed to identify key geometric features of black hole spacetimes—such as the event horizon, the ergosurface—and to provide measures for physical properties like mass and spin Abdelqader and Lake (2015); Tavlayan and Tekin (2020). These scalars are constructed from the Weyl tensor $C_{\mu\nu\alpha\beta}$ and its left dual ${}^{*}C_{\mu\nu\alpha\beta}$, as well as their covariant derivatives. The definitions are as follows:
$\begin{array}[]{ll}I_{1}=C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta},&I_{2}={}^{*}C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta},\\[2.84526pt] I_{3}=\nabla_{\rho}C_{\mu\nu\alpha\beta}\nabla^{\rho}C^{\mu\nu\alpha\beta},&I_{4}=\nabla_{\rho}C_{\mu\nu\alpha\beta}\nabla^{\rho*}C^{\mu\nu\alpha\beta},\\[2.84526pt] I_{5}=k_{\mu}k^{\mu},\quad I_{6}=l_{\mu}l^{\mu},&I_{7}=k_{\mu}l^{\mu}.\end{array}$ (3)
Here, the covectors $k_{\mu}$ and $l_{\mu}$ are defined via the gradients of the first two invariants:
$k_{\mu}=-\nabla_{\mu}I_{1},\quad l_{\mu}=-\nabla_{\mu}I_{2}.$ (4)
These invariants are powerful because, as shown in Tavlayan and Tekin (2020), one may locate the event horizon of Schwarzschild-like black holes by finding where $I_{3}$ vanishes. In these spacetimes, $I_{3}$ is positive outside the horizon, becomes zero precisely at the horizon, and is negative inside.
To illustrate these ideas in a simpler setting, consider the induced metric on a two-dimensional hypersurface with coordinates $(t,r)$. In the induced coordinates, the metric takes the diagonal form
$\gamma_{ij}=\left(\begin{array}[]{cc}f(r)&0\\ 0&\frac{1}{f(r)}.\end{array}\right)$ (5)
where $f(r)$ is the lapse function. This metric form is particularly useful when analyzing horizon properties since it retains the essential radial dependence.
An important quantity computed from the induced metric is the Kretschmann invariant, which for this two-dimensional geometry is given by Pantig and Övgün (2022)
$K={}^{\Sigma}I_{1}={}^{\Sigma}R_{ijkl}{}^{\Sigma}R^{ijkl}=\left({\frac{{\rm d}^{2}}{{\rm d}{r}^{2}}}f(r)\right)^{2},$ (6)
This invariant provides a measure of the curvature intrinsic to the hypersurface. To further probe the horizon, one can define a horizon-detecting invariant based on the gradient of this curvature invariant:
${}^{\Sigma}I_{5}=\nabla_{m}{}^{\Sigma}I_{1}\,\nabla^{m}{}^{\Sigma}I_{1}=4\,\left(\frac{d^{2}}{dr^{2}}f(r)\right)^{2}\left(\frac{d^{3}}{dr^{3}}f(r)\right)^{2}f(r).$ (7)
The prescription is that the largest real root of ${}^{\Sigma}I_{5}$ (i.e. where ${}^{\Sigma}I_{5}(r_{+})=0$) identifies the event horizon. Similarly, one can consider
${}^{\Sigma}I_{3}=\nabla_{m}{}^{\Sigma}R_{ijkl}\,\nabla^{m}{}^{\Sigma}R^{ijkl}=\left(\frac{d^{3}}{dr^{3}}f(r)\right)^{2}f(r),$ (8)
which also vanishes at the horizon. These expressions underline how derivatives of the lapse function encode critical geometric transitions associated with the horizon.
Next, by inserting the lapse function from Eq. (1) into the induced metric Eq. (5), we can compute explicit forms for these invariants. The Kretschmann invariant associated with the induced metric becomes
$K={}^{\Sigma}I_{1}=\frac{m_{S}^{2}}{4\pi^{2}m_{P}^{4}\,r^{6}}.$ (9)
Similarly, the horizon-detecting invariants are found to be
${}^{\Sigma}I_{5}=\frac{9m_{S}^{4}}{4\pi^{4}m_{P}^{8}\,r^{14}}\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right),$ (10)
${}^{\Sigma}I_{3}=\frac{9m_{S}^{2}}{4\pi^{2}m_{P}^{4}\,r^{8}}\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right).$ (11)
The common factor $\left(\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right)$ in these expressions is key, as it directly determines the horizon position. Setting either ${}^{\Sigma}I_{5}$ or ${}^{\Sigma}I_{3}$ to zero yields the event horizon radius
$r_{h}=\frac{A_{0}^{2}\,m_{S}}{4\pi m_{P}^{2}}.$ (12)
Thus, the invariants derived from the induced metric not only detect the presence of the event horizon but also give a clear algebraic relation for its location in terms of the parameters $A_{0}$, $m_{S}$, and $m_{P}$.
Finally, the Kretschmann scalar for the full four-dimensional spacetime is given by
$R^{\alpha\beta\gamma\delta}R_{\alpha\beta\gamma\delta}=\frac{m_{S}^{2}}{2\pi^{2}m_{P}^{4}\,r^{6}}+\frac{4}{r^{4}}\left[1-\frac{1}{A_{0}^{2}}+\frac{m_{S}}{4\pi m_{P}^{2}\,r}\right]^{2}.$ (13)
Overall, these derivations underscore how curvature invariants—constructed from the Weyl tensor and its derivatives—serve as robust tools for detecting horizons and exploring the geometrical properties of black hole spacetimes. They provide an alternative method to the traditional approach of locating horizons by solving $g^{rr}=0$, and offer additional insights into the interplay between the spacetime geometry and the parameters that define the gravitational field.
III Calculation of Hawking Radiation in Jacobi Metric formalism
Recent observational and theoretical advancements have significantly enriched our understanding of black hole physics, particularly concerning intriguing features such as black hole shadows, photon spheres, and gravitational lensing. Hawking radiation, conceptualized via tunneling mechanisms, has provided profound insights into quantum aspects of black holes Akhmedov et al. (2006); Parikh and Wilczek (2000); Shankaranarayanan et al. (2002); Angheben et al. (2005). To compute the Hawking temperature of a black hole, we adopt a semi-classical tunneling approach that utilizes the Jacobi metric derived from the full four-dimensional covariant metric Bera et al. (2020). In this framework, the tunneling probability for a particle crossing the event horizon is related to the imaginary part of the classical action Akhmedov et al. (2006); Kerner and Mann (2008a, 2006, b); Yale and Mann (2009). We begin by considering the particle’s wave function in the WKB approximation, written as
$\psi=e^{\frac{i}{\hbar}S},$ (14)
where S is the classical action. The key idea is that the tunneling probability is dominated by the exponential of the imaginary part of S.
The corresponding Jacobi metric, which encapsulates the effective geometry experienced by the particle, is given by Gibbons (2016); Chanda et al. (2017); Das et al. (2017):
$\mathrm{d}s^{2}=j_{ij},\mathrm{d}x^{i}\mathrm{d}x^{j}$
$=\Big{(}E^{2}-m^{2}f(r)\Big{)}\left(\frac{\mathrm{d}r^{2}}{f^{2}(r)}+\frac{r^{2}}{f(r)}\Big{(}\mathrm{d}\theta^{2}+\sin^{2}\theta,\mathrm{d}\phi^{2}\Big{)}\right).$ (15)
Here, E and m are the energy and mass of the tunneling particle, respectively, and $f(r)$ is the lapse function that encodes the gravitational potential. This metric effectively redefines the “distance” in configuration space in such a way that the kinetic term of the particle’s motion is modulated by the combination $E^{2}-m^{2}f(r)$.
In the context of Hawking radiation, the tunneling process occurs predominantly along the radial direction near the horizon. Thus, the action for the particle undergoing radial tunneling can be written as
$S=-\int\sqrt{j_{ij},\frac{\mathrm{d}x^{i}}{\mathrm{d}s},\frac{\mathrm{d}x^{j}}{\mathrm{d}s}},\mathrm{d}s~{}.$ (16)
The integrand represents the proper “speed” of the particle in the effective geometry. Explicitly, for the radial component, this becomes
$\sqrt{j_{ij},\frac{\mathrm{d}x^{i}}{\mathrm{d}s},\frac{\mathrm{d}x^{j}}{\mathrm{d}s}}=\pm\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r)\frac{\mathrm{d}r}{\mathrm{d}s}~{}.$ (17)
The choice of the $\pm$ sign distinguishes between incoming and outgoing trajectories. From Eq. (17) and the definition of the action in Eq. (16), the radial momentum of the particle is obtained as Bera et al. (2020)
$p_{r}=\partial_{r}S=\mp\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r)~{}.$ (18)
Since the tunneling process is confined to the near-horizon region, the condition $f(r)<0$ is imposed. In our expression for $p_{r}$, the sign ensures that the outgoing particle (with $p_{r}>0$) corresponds to the appropriate tunneling direction, in line with conventional treatments such as those found in Srinivasan and Padmanabhan (1999).
Near the horizon, the full four-dimensional metric simplifies to an effective (1+1)-dimensional form because the angular components become subdominant Robinson and Wilczek (2005); Iso et al. (2006); Majhi (2010). To capture the physics near the horizon, we expand the lapse function $f(r)$ in a Taylor series about the horizon $r=r_{H}$:
$f(r)=f(r_{H})+f’(r_{H})(r-r_{H})+\mathcal{O}((r-r_{H})^{2})$
$\approx 2\tilde{\kappa}(r-r_{H})+\mathcal{O}((r-r_{H})^{2})~{},$ (19)
where we have used the fact that $f(r_{H})=0$ and defined the surface gravity as
$\tilde{\kappa}=\frac{1}{2}f’(r_{H})~{}.$ (20)
Substituting the expansion (III) into Eq. (17), the near-horizon radial action becomes Bera et al. (2020)
$S=\mp\frac{E}{2\tilde{\kappa}}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}\pm\frac{m^{2}}{2E}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathrm{d}r$
$\mp\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathcal{O}(r-r_{H}),\mathrm{d}r~{}$ (21)
where $\epsilon$ is a small positive parameter that spans a region across the horizon. In the first integral, we perform a change of variable $r-r_{H}=\epsilon e^{i\theta}$ to evaluate the contour integral. This yields
$\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}=-i\pi~{}.$ (22)
The second integral vanishes in the limit $\epsilon\to 0$, and the contributions from higher-order terms (third integral) are negligible. Thus, the dominant contribution to the action is
[…]
Here, the positive sign corresponds to the outgoing trajectory, while the negative sign corresponds to the incoming trajectory.
The WKB wave functions for the outgoing and incoming particles can now be written as
$\psi_{\rm out}=A,e^{\frac{i}{\hbar}S_{\rm out}},\quad\psi_{\rm in}=A,e^{\frac{i}{\hbar}S_{\rm in}},$ (24)
with A being a normalization constant. The associated probabilities are then
$P_{\rm out}=|\psi_{\rm out}|^{2}=|A|^{2}e^{-\frac{\pi E}{\hbar\tilde{\kappa}}},$ (25)
and
$P_{\rm in}=|\psi_{\rm in}|^{2}=|A|^{2}e^{\frac{\pi E}{\hbar\tilde{\kappa}}}~{}.$ (26)
Since the real part of the action does not contribute to the probability, the tunneling rate is determined by the ratio
$\Gamma=\frac{P_{\rm out}}{P_{\rm in}}=e^{-\frac{2\pi E}{\hbar\tilde{\kappa}}}\equiv e^{-\frac{E}{T_{H}}}~{}.$ (27)
Comparing Eq. (27) with the Boltzmann factor, we identify the Hawking temperature as
$T_{H}=\frac{\tilde{\kappa}}{2\pi}~{}.$ (28)
To further clarify, the surface gravity $\tilde{\kappa}$ can also be expressed in terms of the norm of the timelike Killing vector field $\chi^{\mu}$ as
$\tilde{\kappa}=\left.\sqrt{-\frac{1}{2}\nabla_{\mu}\chi_{\nu}\nabla^{\mu}\chi^{\nu}},\right|{r=r_{H}}=\left.\frac{1}{2}\frac{\partial f(r)}{\partial r},\right|{r=r_{H}}~{}.$ (29)
In our model, using the explicit form of the lapse function, the Hawking temperature can be further determined as
$T_{H}=\frac{m_{p}^{2}}{A_{0}^{4}m_{s}}~{},$ (30)
which encapsulates the dependence on the parameters $m_{p}$, $m_{s}$, and the scale parameter $A_{0}$.
Figure 1 illustrates how the Hawking temperature $T_{H}$ varies with the black hole mass parameter $m_{s}$ for different values of $A_{0}$. Figure 1 shows that as $m_{s}$ increases, the temperature rapidly decreases, indicating that heavier black holes radiate less intensely. Additionally, smaller values of $A_{0}$ result in higher temperatures, implying that configurations with lower $A_{0}$ are thermodynamically less stable.
[…]
In summary, this semi-classical tunneling method based on the Jacobi metric provides a consistent framework to derive the Hawking temperature. The detailed derivation—from the form of the Jacobi metric through the evaluation of the tunneling action and the resulting probability—shows that the Hawking temperature is intimately linked to the surface gravity of the black hole. This method not only reinforces the interpretation of Hawking radiation as a quantum tunneling process but also highlights the interplay between the black hole’s geometric parameters and its thermodynamic properties.
IV Weak Deflection Angle Using the Gauss-Bonnet Theorem
In this section, we analyze the weak gravitational lensing phenomena caused by the black hole using the Gauss–Bonnet theorem Gibbons and Werner (2008). This approach involves studying the deflection of photons traveling in the spacetime geometry induced by the black hole. To start, we consider the optical metric projected onto the equatorial plane defined by $\theta=\pi/2$. This optical geometry is represented by the metric:
$\mathrm{d}t^{2}=\frac{1}{f(r)^{2}}\mathrm{d}r^{2}+\frac{r^{2}}{f(r)}\mathrm{d}\varphi^{2}.$ (31)
The Gaussian optical curvature associated with this metric, which measures the intrinsic curvature affecting photon trajectories, is calculated using the Ricci scalar R as a relation $\mathcal{K}=R/2$. Explicitly, Gaussian curvature becomes:
$\mathcal{K}=\frac{3m_{s}^{2}}{64\pi^{2}m_{p}^{4}r^{4}}-\frac{m_{s}}{4\pi A_{0}^{2}m_{p}^{2}r^{3}}.$ (32)
To derive the weak deflection angle using the Gauss–Bonnet theorem, we first define a non-singular region $\mathcal{D}_{R}$ bounded by the curve $\partial\mathcal{D}_{R}=\gamma_{\tilde{g}}\cup C_{R}$. Applying the Gauss–Bonnet theorem yields Gibbons and Werner (2008):
$\iint\limits_{\mathcal{D}_{R}}\mathcal{K}\,\mathrm{d}S+\oint\limits_{\partial\mathcal{D}_{R}}\kappa\,\mathrm{d}t+\sum_{i}\theta_{i}=2\pi\chi(\mathcal{D}_{R}),$ (33)
where $\kappa$ represents the geodesic curvature of the boundary, $\theta_{i}$ is the exterior angle at the $i^{th}$ vertex, and $\chi(\mathcal{D}_{R})=1$ denotes the Euler characteristic for a simply connected domain.
We choose our region outside the photon trajectory, and since the photon path $\gamma_{\tilde{g}}$ is geodesic, its geodesic curvature vanishes, i.e., $\kappa(\gamma_{\tilde{g}})=0$. For the boundary at infinity, represented by the curve $C_{R}:=r(\varphi)=R=\text{constant}$, the geodesic curvature is computed via:
$\kappa(C_{R})=|\nabla_{\dot{C}_{R}}\dot{C}_{R}|.$ (34)
Expanding this expression further, we have the radial component:
$(\nabla_{\dot{C}_{R}}\dot{C}_{R})^{r}=\dot{C}_{R}^{\varphi}(\partial_{\varphi}\dot{C}_{R}^{r})+\tilde{\Gamma}_{\varphi\varphi}^{r}(\dot{C}_{R}^{\varphi})^{2}.$ (35)
The first term in Eq. (35) vanishes due to the constancy of $R$. Evaluating the second term using the unit-speed condition, we obtain:
$\lim_{R\rightarrow\infty}\kappa(C_{R})=\frac{1}{R}.$ (36)
At large radial distances, we further approximate:
$\lim_{R\rightarrow\infty}\mathrm{d}t\approx R\,\mathrm{d}\varphi,$ (37)
thus combining Eqs. (36) and (37), we have $\kappa(C_{R})\mathrm{d}t=\mathrm{d}\varphi$.
Using the straight-line approximation for photon trajectories, the radial coordinate is related to the impact parameter $b$ as $r=b/\sin\varphi$. Consequently, the Gauss–Bonnet theorem simplifies considerably, providing a straightforward integral representation for the deflection angle Gibbons and Werner (2008):
$\alpha=-\int_{0}^{\pi}\int_{\frac{b}{\sin\varphi}}^{\infty}\mathcal{K}\,\mathrm{d}S,$ (38)
with the surface element given explicitly as:
$\mathrm{d}S=\left(\frac{3A_{0}^{5}m_{s}}{8\pi m_{p}^{2}}+A_{0}^{3}r\right)\mathrm{d}r\,\mathrm{d}\varphi.$ (39)
Evaluating the integral (38) using the Gaussian curvature from Eq. (32), we find the weak deflection angle, including second-order contributions:
$\alpha\approx\frac{3A_{0}^{3}m_{s}^{2}}{256\pi b^{2}m_{p}^{4}}+\frac{A_{0}m_{s}}{2\pi bm_{p}^{2}}.$ (40)
Figure 2: Weak deflection angle plotted as a function of radial distance parameter $r$ for various values of the scale parameter $A_{0}$. The Planck mass squared is fixed at $m_{p}^{2}=1$.
Figure 2 illustrates how the weak deflection angle depends on the impact parameter $b$ for different values of the parameter $A_{0}$. Notably, the deflection angle decreases gradually as one moves away from the black hole, underscoring how gravitational lensing diminishes at larger distances. Furthermore, varying $A_{0}$ significantly influences lensing strength, with smaller values yielding stronger gravitational deflection.
IV.1 Derivation of deflection angle using geodesics method
We now independently verify the deflection angle by solving the geodesic equations directly. Introducing the impact parameter $b\equiv L/E$, we write
$\dot{t}=\frac{E}{f(r)},\quad\dot{\phi}=\frac{L}{r^{2}}\,,$ (41)
which follow from symmetries of the spacetime and corresponding constants of motion $E$ and $L$.
Substituting these expressions into the null geodesic condition ($\mathcal{L}=0$), we have
$-f(r)\left(\frac{E}{f(r)}\right)^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\left(\frac{L}{r^{2}}\right)^{2}=0\,,$ (42)
which simplifies to
$\dot{r}^{2}=E^{2}-\frac{L^{2}}{r^{2}}\,f(r)\,.$ (43)
The above equation is the radial geodesic equation, expressing the radial velocity in terms of the energy and angular momentum constants of motion.
Introducing the inverse radial coordinate $u\equiv 1/r$, we note
$\frac{dr}{d\phi}=-\frac{1}{u^{2}}\frac{du}{d\phi}\,,\quad\dot{\phi}=Lu^{2}\,,$ (44)
thus giving
$\dot{r}=\frac{dr}{d\phi}\dot{\phi}=-L\frac{du}{d\phi}\,.$ (45)
Substituting this into the radial equation yields
$L^{2}\left(\frac{du}{d\phi}\right)^{2}=E^{2}-L^{2}u^{2}\,f\left(\frac{1}{u}\right)\,.$ (46)
Dividing by $L^{2}$ and using the impact parameter $b$, we find
$\left(\frac{du}{d\phi}\right)^{2}=\frac{1}{b^{2}}-u^{2}\,f\left(\frac{1}{u}\right)\,.$ (47)
Expanding the metric function $f(1/u)$, we have explicitly
$f\left(\frac{1}{u}\right)=\frac{1}{A_{0}^{2}}-\frac{m_{S}}{4\pi m_{P}^{2}}\,u\,,$ (48)
thus yielding the orbital equation
$\left(\frac{du}{d\phi}\right)^{2}=\frac{1}{b^{2}}-\frac{u^{2}}{A_{0}^{2}}+\frac{m_{S}}{4\pi m_{P}^{2}}u^{3}\,.$ (49)
This equation explicitly relates the inverse radius $u$ and the azimuthal angle $\phi$.
Differentiating Eq. (49) again gives the second-order orbital equation:
$\frac{d^{2}u}{d\phi^{2}}+\frac{u}{A_{0}^{2}}=\frac{3m_{S}}{8\pi m_{P}^{2}}u^{2}\,,$ (50)
which governs the photon trajectory around the massive source.
To solve perturbatively, let $u(\phi)=u_{0}(\phi)+\delta u(\phi)$, where $u_{0}$ solves the unperturbed equation (with $m_{S}=0$) and $\delta u$ is a small correction.
Unperturbed solution:
For $m_{S}=0$, Eq. (50) reduces to a harmonic oscillator-type equation
$\frac{d^{2}u_{0}}{d\phi^{2}}+\frac{u_{0}}{A_{0}^{2}}=0\,,$ (51)
whose physically relevant solution is
$u_{0}(\phi)=\frac{1}{b}\sin\left(\frac{\phi}{A_{0}}\right)\,.$ (52)
This solution describes a straight trajectory (no deflection), and the closest approach is at $\phi=\pi A_{0}/2$.
First-order correction:
The perturbation satisfies
$\frac{d^{2}\,\delta u}{d\phi^{2}}+\frac{\delta u}{A_{0}^{2}}=\frac{3m_{S}}{8\pi m_{P}^{2}}u_{0}^{2}(\phi)\,.$ (53)
Expanding $u_{0}^{2}(\phi)$ explicitly, we have
$u_{0}^{2}(\phi)=\frac{1}{2b^{2}}\left[1-\cos\left(\frac{2\phi}{A_{0}}\right)\right]\,,$ (54)
thus yielding
$\frac{d^{2}\,\delta u}{d\phi^{2}}+\frac{\delta u}{A_{0}^{2}}=\frac{3m_{S}}{16\pi m_{P}^{2}b^{2}}\left[1-\cos\left(\frac{2\phi}{A_{0}}\right)\right]\,.$ (55)
We try a particular solution of the form
$\delta u_{p}(\phi)=C+D\cos\left(\frac{2\phi}{A_{0}}\right)\,.$ (56)
Substituting and matching coefficients yields
$C=\frac{3m_{S}A_{0}^{2}}{16\pi m_{P}^{2}b^{2}}\,,\quad D=\frac{m_{S}A_{0}^{2}}{16\pi m_{P}^{2}b^{2}}\,.$ (57)
Thus, the particular solution is
$\delta u_{p}(\phi)=\frac{3m_{S}A_{0}^{2}}{16\pi m_{P}^{2}b^{2}}+\frac{m_{S}A_{0}^{2}}{16\pi m_{P}^{2}b^{2}}\cos\left(\frac{2\phi}{A_{0}}\right)\,.$ (58)
Adding suitable homogeneous solutions and applying physical boundary conditions (that the orbit is asymptotically straight), we obtain the full corrected trajectory. Finally, by examining the asymptotic limit $u\to 0$, the total deflection angle $\alpha$ can be extracted. Performing this carefully (as in standard methods from the literature Dolan (2023)), the final deflection angle up to first order in perturbation is obtained:
$\alpha\approx\frac{3A_{0}^{3}m_{S}^{2}}{256\pi b^{2}m_{P}^{4}}+\frac{A_{0}m_{S}}{2\pi bm_{P}^{2}}\,.$ (59)
The first term represents a second-order correction, while the second term is the leading-order deflection angle, reproducing known weak-field lensing results.
V Shadow of the Black Hole
The photon sphere around a black hole is defined by unstable circular photon orbits that are significantly influenced by the gravitational field. Photons in these orbits either spiral into the black hole’s event horizon or escape to infinity, forming a luminous photon ring. This ring delineates the observable boundary of the black hole shadow for distant observers Synge (1966); Cunningham and Bardeen (1972); Bardeen et al. (1972); Luminet (1979a). Historically, different terminologies such as escape cone Synge (1966), cone of gravitational radiation capture Zeldovich and Novikov (1971), optical appearance of black holes, and black hole image Bardeen (1973); Chandrasekhar (1998); Luminet (1979b, 2018) have been utilized. However, the contemporary and widely adopted term black hole shadow, introduced by Falcke et al. Falcke et al. (2000), refers explicitly to the dark region surrounded by the photon ring Johannsen and Psaltis (2010c); Johnson et al. (2020).
In this section, we investigate the properties of the black hole shadow, focusing particularly on its dependence on the underlying model parameters, such as $A_{0}$, $m_{s}$, and $m_{p}$. Given the spherical symmetry of the spacetime considered, the analysis is simplified by restricting null geodesic trajectories to the equatorial plane ($\theta=\pi/2$). The metric in this case reduces to:
$\mathrm{d}s^{2}=-A(r)\mathrm{d}t^{2}+B(r)\mathrm{d}r^{2}+C(r)\mathrm{d}\phi^{2},$ (60)
where $B(r)=A(r)^{-1}$ and $C(r)=r^{2}$. The geodesic motion of photons can be described through the Lagrangian:
$2\mathcal{L}=-A(r)\dot{t}^{2}+B(r)\dot{r}^{2}+C(r)\dot{\phi}^{2},$ (61)
where the dot indicates differentiation with respect to an affine parameter $\lambda$. Constants of motion, namely the energy $E$ and angular momentum $L$, arise from symmetry: $E=A(r)\dot{t},\quad L=C(r)\dot{\phi}.$ Using these constants, the critical impact parameter $b$, which quantifies the photon trajectory, is defined as:
$b=\frac{L}{E}=\frac{C(r)}{A(r)}\frac{\mathrm{d}\phi}{\mathrm{d}t}.$ (62)
Applying the null condition $ds^{2}=0$, the radial equation for photon trajectories becomes:
$\left(\frac{\mathrm{d}r}{\mathrm{d}\phi}\right)^{2}=\frac{C(r)}{B(r)}\left[\frac{h(r)^{2}}{b^{2}}-1\right],$ (63)
where
$h(r)^{2}=\frac{C(r)}{A(r)}.$ (64)
The photon sphere is determined by the condition $\frac{\mathrm{d}}{\mathrm{d}r}h^{2}(r)=0$, explicitly:
$C^{\prime}(r_{\mathrm{ph}})A(r_{\mathrm{ph}})-C(r_{\mathrm{ph}})A^{\prime}(r_{\mathrm{ph}})=0,$ (65)
yielding the photon sphere radius:
$r_{\mathrm{ph}}=\frac{3A_{0}^{2}m_{S}}{8\pi m_{P}^{2}}.$ (66)
The black hole shadow radius $r_{\mathrm{sh}}$, as seen by an observer at radial coordinate $r_{O}$, is:
$r_{\mathrm{sh}}=r_{O}\sin\beta=r_{\mathrm{ph}}\sqrt{\frac{1}{A(r_{\mathrm{ph}})}},$ (67)
leading explicitly to:
$\boxed{r_{\mathrm{sh}}=\frac{3\sqrt{3}A_{0}^{3}m_{s}}{8\pi m_{p}^{2}}}.$ (68)
Considering a Schwarzschild black hole scenario (with $m_{s}=1$) and a known shadow radius $r_{\mathrm{sh}}=5.2$, we derive the parameter relationship:
$\frac{A_{0}^{3}}{m_{p}^{2}}=\frac{8\pi r_{\mathrm{sh}}}{3\sqrt{3}}\approx 25.164.$ (69)
This relation constrains the combination $A_{0}^{3}/m_{p}^{2}$, yet it allows flexibility in selecting either parameter individually. For example:
$A_{0}\approx 2.93\quad\text{if}\quad m_{p}=1,$ (70)
$m_{p}\approx 0.199\quad\text{if}\quad A_{0}=1.$ (71)
Figures 3 and 4 highlight the sensitivity of the photon sphere and shadow radii to the parameter $A_{0}$. Both radii exhibit monotonic growth with increasing $A_{0}$. The shadow radius, however, demonstrates a notably stronger dependence due to its cubic relationship with $A_{0}$, indicating how alterations in $A_{0}$ significantly affect the geometry around the black hole. This result underscores the potential observational impact of different model parameters in gravitational lensing phenomena.
Figure 3: Shadow radius and photon sphere radius as functions of the parameter $A_{0}$, assuming $m_{p}=m_{s}=1$.
Figure 4: Visualization of the black hole shadow with an accretion disk, highlighting the geometrical effect of parameter variations ($m_{p}=m_{s}=1$).
VI Strong lensing by the black hole
Recalling the form in Eq. (61) for null geodesics (i.e., with $\mathcal{L}=0$), we assume that light rays approach the black hole at the radial distance $r=r_{0}$. This simplifies the trajectory equation at this turning point to
$A_{0}\dot{t}_{0}^{2}=C_{0}\dot{\phi}_{0}^{2},$ (72)
where $X_{0}\equiv X(r_{0})$. Using the definition of the impact parameter in Eq. (62) at $r=r_{0}$, the function $R(r)$ in Eq. (63) can be rewritten as
$R(r)=\frac{A_{0}C}{AC_{0}}-1.$ (73)
We define the critical impact parameter as Bozza (2002b)
$b_{c}\equiv b(r_{\mathrm{ph}})=\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}\sqrt{\frac{C_{0}}{A_{0}}}.$ (74)
In Ref. Bozza (2002b), the deflection angle of light rays passing at the radial distance $r_{0}$ from a static asymptotically flat black hole is given as
[…]
where $R(r)$ is given in Eq. (73). Introducing the variable $z\equiv 1-r_{0}/r$, the integral in Eq. (75) can be separated into the divergent part, $I_{D}(r_{0})$, and the regular part, $I_{R}(r_{0})$. The divergent part is expressed as Bozza (2002b); Tsukamoto (2018); Ali et al. (2025)
[…]
where
$f_{0}(z,r_{0})=\frac{2r_{0}}{\sqrt{x_{1}(r_{0})z+x_{2}(r_{0})z^{2}}},$ (77)
and
$x_{1}(r_{0})=\frac{1-A_{0}}{C_{0}A_{0}^{\prime}}\left(C_{0}^{\prime}A_{0}-C_{0}A_{0}^{\prime}\right),$ (78a)
$x_{2}(r_{0})=\frac{(1-A_{0})^{2}}{2C_{0}^{2}A_{0}^{\prime 3}}\Big{[}2C_{0}C_{0}^{\prime}A_{0}^{\prime 2}+(C_{0}C_{0}^{\prime\prime}-2C_{0}^{\prime 2})A_{0}A_{0}^{\prime}-C_{0}C_{0}^{\prime}A_{0}A_{0}^{\prime\prime}\Big{]}.$ (78b)
This integral can be evaluated analytically, yielding
$I_{D}(r_{0})=\frac{4r_{0}}{\sqrt{x_{2}(r_{0})}}\ln\left(\frac{\sqrt{x_{2}(r_{0})}+\sqrt{x_{1}(r_{0})+x_{2}(r_{0})}}{\sqrt{x_{1}(r_{0})}}\right).$ (79)
In the strong lensing regime, i.e., when $r_{0}\rightarrow r_{\mathrm{ph}}$, expanding $x_{1}(r_{0})$ around $r_{0}-r_{\mathrm{ph}}$ gives
$x_{1}(r_{0})=\frac{C_{\mathrm{ph}}r_{\mathrm{ph}}\mathfrak{L^{\prime}_{\mathrm{ph}}}}{B_{\mathrm{ph}}}\left(r_{0}-r_{\mathrm{ph}}\right)+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2},$ (80)
where $\mathfrak{L}(r)=C^{\prime}(r)/C(r)-A^{\prime}(r)/A(r)$. The impact parameter is similarly expanded as
$b(r_{0})=b_{c}+\frac{1}{4}\sqrt{\frac{C_{\mathrm{ph}}}{A_{\mathrm{ph}}}}\,\mathfrak{L}_{\mathrm{ph}}^{\prime}\left(r_{0}-r_{\mathrm{ph}}\right)^{2}+\mathcal{O}(r_{0}-r_{\mathrm{ph}})^{2}.$ (81)
Thus, we obtain
$\lim_{r_{0}\rightarrow r_{\mathrm{ph}}}x_{1}(r_{0})=\lim_{b\rightarrow b_{c}}\frac{2C_{\mathrm{ph}}r_{\mathrm{ph}}\sqrt{\mathfrak{L}_{\mathrm{ph}}^{\prime}}}{B_{\mathrm{ph}}}\left(\frac{b}{b_{c}}-1\right)^{1/2}.$ (82)
This provides
$I_{D}(b)=-\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\left(\frac{b}{b_{c}}-1\right)+\frac{r_{\mathrm{ph}}}{\sqrt{x_{2}(r_{\mathrm{ph}})}}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}^{\prime}_{\mathrm{ph}}\Bigr{)}+\mathcal{O}(b-b_{c}),$ (83)
where
$x_{2}(r_{\mathrm{ph}})=\frac{C_{\mathrm{ph}}(1-A_{\mathrm{ph}})^{2}\Bigl{[}C_{\mathrm{ph}}^{\prime\prime}A_{\mathrm{ph}}-C_{\mathrm{ph}}A^{\prime\prime}_{\mathrm{ph}}\Bigr{]}}{2A_{\mathrm{ph}}^{2}C_{\mathrm{ph}}^{\prime}}.$ (84)
The regular part of the integral in Eq. (75), $I_{R}$, is expressed as
[…]
Thus, the deflection angle in the strong lensing regime is given by
$\alpha(b)=-\bar{a}\ln\left(\frac{b}{b_{c}}-1\right)+\bar{k}+\mathcal{O}(b-b_{c}),$ (86)
where
$\bar{a}=\sqrt{\frac{2B_{\mathrm{ph}}A_{\mathrm{ph}}}{C^{\prime\prime}_{\mathrm{ph}}A_{\mathrm{ph}}-C_{\mathrm{ph}}A^{\prime\prime}_{\mathrm{ph}}}},$ (87a)
$\bar{k}=\bar{a}\ln\Bigl{(}r_{\mathrm{ph}}^{2}\mathfrak{L}_{\mathrm{ph}}^{\prime}\Bigr{)}+I_{R}(r_{\mathrm{ph}})-\pi.$ (87b)
Now, using the spacetime metric functions from the line element (2) for the black hole in LGT and comparing them with those in Eq. (60), while considering the photon sphere radius given in Eq. (66), we obtain the critical impact parameter as
$b_{c}=\sqrt{\frac{C_{\mathrm{ph}}}{A_{\mathrm{ph}}}}=\frac{3\sqrt{3}A_{0}^{3}m_{S}}{8\pi m_{P}^{2}},$ (88)
which coincides with the shadow radius presented in Eq. (68). By utilizing the relations derived in this section, we further obtain
$\bar{a}=A_{0},$ (89)
$\bar{k}=A_{0}\ln\left(2+\frac{256\pi^{2}m_{P}^{4}}{27A_{0}^{6}m_{S}^{2}}\right)+I_{R}(r_{\mathrm{ph}})-\pi.$ (90)
The integrand of Eq. (85) is given by
$\frac{2\left(3A_{0}^{2}-1\right)\left[\sqrt{-18+\frac{3}{A_{0}^{2}}+9A_{0}^{2}}-\sqrt{-\frac{\left(3A_{0}^{2}-1\right)^{2}\left(3A_{0}^{2}z-3-z\right)}{A_{0}^{2}}}\right]}{\sqrt{\frac{\left(3A_{0}^{2}-1\right)^{4}\left[3-\left(3A_{0}^{2}-1\right)z\right]}{A_{0}^{4}}}}.$ (91)
Since this integral cannot be computed analytically for general $A_{0}$, we note that for the Schwarzschild black hole case ($A_{0}=\pm 1$), the integral evaluates to $I_{R}^{\mathrm{SBH}}\approx 0.9496$, which matches the result reported in Ref. Bozza (2002b). More generally, as $A_{0}^{2}\to 1$, the values of $I_{R}$ approach this Schwarzschild black hole value (see Fig. 5).
Figure 5: The behavior of $I_{R}(r_{\mathrm{ph}})$ as a function of $A_{0}$. The dashed line corresponds to $I_{R}=0.9496$, which is the value for the Schwarzschild black hole lensing.
Additionally, for $A_{0}^{2}=1$, we find $\bar{k}=-0.40023$, which corresponds to the Schwarzschild black hole lensing result reported in Ref. Bozza (2002b). Finally, by incorporating these relations, we present in Fig. 6 the $b$-profile of the strong deflection angle as a function of $A_{0}$.
Figure 6: The $b$-profile of the strong deflection angle $\alpha$, plotted for $0<A_{0}\leq 1$.
VI.1 The lens equation and the EHT constraints
Following the discussion in Ref. Bozza et al. (2001), we consider a scenario in which a light source, denoted as $S$, is nearly perfectly aligned with a black hole that serves as a gravitational lens, labeled as $L$. In this configuration, relativistic images are expected to form. The lens equation describing this setup is given by
$\psi=\theta-\frac{D_{LS}}{D_{OS}}\Delta\alpha_{n},$ (92)
where $D_{LS}$ represents the distance between the lens and the light source, while $D_{OS}$ denotes the distance between the observer and the light source. The angular positions of the source and the image are given by $\psi$ and $\theta$, respectively. Additionally, the quantity $\Delta\alpha_{n}=\alpha(\theta)-2n\pi$ accounts for the deviation of the deflection angle after considering the total number of loops completed by the photons. For $\alpha(\theta_{n}^{0})=2n\pi$, we obtain
$\theta_{n}^{0}=\frac{b_{c}}{D_{OL}}\left(1+\epsilon_{n}\right),$ (93)
where
$\epsilon_{n}=\exp\left(\frac{\bar{k}-2n\pi}{\bar{a}}\right).$ (94)
Expanding $\alpha(\theta)$ around $\theta_{n}^{0}$ and introducing $\Delta\theta_{n}=\theta-\theta_{n}^{0}$, the deviation simplifies to
$\Delta\alpha_{n}=-\frac{\bar{a}D_{OL}}{b_{c}\epsilon_{n}}\Delta\theta_{n}.$ (95)
Thus, the lens equation can be rewritten as
$\psi=\theta+\left(\frac{\bar{a}D_{LS}D_{OL}}{b_{c}\epsilon_{n}D_{OS}}\right)\Delta\theta_{n}.$ (96)
Assuming that $D_{OL}\gg b_{c}$, the angular position of the $n$th relativistic image is given by Bozza (2002c)
$\theta_{n}=\theta_{n}^{0}+\frac{b_{c}\epsilon_{n}\left(\psi-\theta_{n}^{0}\right)D_{OS}}{\bar{a}D_{LS}D_{OL}}.$ (97)
From this equation, it follows that when $\psi=\theta_{n}^{0}$, the image coincides with the source. The sign of $\psi$ determines whether the image forms on the same side ($\psi>0$) or the opposite side ($\psi<0$) of the lens. In cases where the black hole, source, and observer are nearly aligned ($\psi\approx 0$), and assuming that the observer and lens are equidistant from the source ($D_{OS}=D_{LS}=2D_{OL}$), light deflection occurs symmetrically, leading to the formation of Einstein rings Einstein (1936); Mellier (1999); Bartelmann and Schneider (2001); Schmidt (2008); Guzik et al. (2010). Under these conditions, Eq. (97) simplifies to Bozza and Mancini (2004)
$\theta_{n}^{E}=\left(1-\frac{b_{c}\epsilon_{n}D_{OS}}{\bar{a}D_{LS}D_{OL}}\right)\theta_{n}^{0}.$ (98)
Furthermore, for $D_{OL}\gg b_{c}$, the angular radius of the $n$th relativistic Einstein ring is given by
$\theta_{n}^{E}=\frac{b_{c}\left(1+\epsilon_{n}\right)}{D_{OL}}.$ (99)
To apply these theoretical results to astrophysical observations, we consider the supermassive black holes M87* and Sgr A*. The black hole M87* has a mass of $(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$ and is located at a distance of $D_{OL}=16.8\,\mathrm{Mpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2019a, b). Meanwhile, Sgr A* has a mass of $4^{+1.1}_{-0.6}\times 10^{6}\,M_{\odot}$ and is situated at a distance of $7.97\,\mathrm{kpc}$ from Earth [The Event Horizon Telescope Collaboration] et al. (2022a); Akiyama et al. (2022). Substituting these parameters into Eq. (99), Fig. 7 illustrates the outermost Einstein rings ($n=1$) for M87* and Sgr A*, assuming they are black holes within the framework of the LGT. These rings are depicted in the celestial coordinate system of an observer on Earth, with coordinates $X$ and $Y$, and are computed for different values of $A_{0}$.
(a) Refer to caption (b)
Figure 7: The Einstein rings of M87* and Sgr A* considered as LGT black holes.
As observed, an increase in $A_{0}$ leads to a significant enlargement of the Einstein rings. Consequently, the Schwarzschild black hole (SBH) exhibits the largest rings when $A_{0}^{2}\leq 1$.
A key quantity in strong lensing is the magnification of the $n$th relativistic image, defined as Virbhadra and Ellis (2000c); Bozza (2002c)
$\mu_{n}=\left.\left(\frac{\psi}{\theta}\frac{d\psi}{d\theta}\right)^{-1}\right|_{\theta_{n}^{0}}=\frac{b_{c}^{2}\epsilon_{n}\left(1+\epsilon_{n}\right)D_{OS}}{\bar{a}\psi D_{LS}D_{OL}^{2}}.$ (100)
Since the magnification is inversely proportional to $D_{OL}^{2}$, relativistic images are typically faint. The outermost image appears the brightest, while the brightness of higher-order images diminishes exponentially. However, in the limiting case where $\psi$ approaches zero, implying near-perfect alignment of the source and the lens, the images undergo significant magnification. It is noteworthy that while the outermost image, corresponding to $\theta_{1}$, remains distinct, higher-order images tend to cluster around $\theta_{\infty}\equiv\theta_{n}|_{n\to\infty}$ Bozza (2002c), which is identified as $\theta_{\infty}=\theta_{c}$, representing the innermost relativistic image.
Astrophysical applications of strong lensing also involve two important observables. The first is the angular separation between the outermost and innermost relativistic images, given by
$s=\theta_{1}-\theta_{\infty}\approx\theta_{\infty}\epsilon_{1},$ (101)
while the second is the relative magnification between the outermost relativistic image and the collective group of inner relativistic images, expressed as Bozza (2002c)
$r_{\mathrm{mag}}=\frac{\mu_{1}}{\sum_{n=2}^{\infty}\mu_{n}}=2.5\log_{10}\left(\exp\left[\frac{2\pi}{\bar{a}}\right]\right),$ (102)
which notably remains independent of the observer’s distance, $D_{OL}$.
Before concluding this section, it is crucial to establish constraints on the scalar hair parameter of black holes based on recent observations of M87* and Sgr A*.
In 2019, the Event Horizon Telescope (EHT) collaboration published the first horizon-scale image of the supermassive black hole M87*, providing strong observational evidence for black holes. Their analysis revealed that the compact emission region had an angular diameter of $\theta_{d}=(42\pm 3)\,\mu\text{as}$, along with a central flux depression exceeding a factor of $\gtrsim 10$, corresponding to the black hole’s shadow [The Event Horizon Telescope Collaboration] et al. (2019c, a, b). Subsequently, in 2022, the EHT collaboration captured an image of Sgr A*, the supermassive black hole at the center of the Milky Way. This image displayed a distinctive ring structure with an angular diameter of $\theta_{d}=(48.7\pm 7)\,\mu\text{as}$ and a deviation from the Schwarzschild shadow characterized by $\delta=-0.08^{+0.09}_{-0.09}$ (VLTI) and $\delta=-0.04^{+0.09}_{-0.10}$ (Keck). Additionally, EHT observations of Sgr A* estimated the angular diameter of the emission ring to be $\theta_{d}=(51.8\pm 2.3)\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022a, b).
Using these EHT observational data, we model M87* and Sgr A* as LGT black holes to constrain the parameter $A_{0}$. By considering the apparent radius of the photon sphere, $\theta_{\infty}$, as the angular size of the black hole shadow, we derive constraints on $A_{0}$ at the $1\sigma$ confidence level. To establish this relationship, we use $\theta_{d}=2\theta_{\infty}$, linking the observed angular diameter of the shadows to the theoretical prediction. In Fig. 8(a), we present the $A_{0}$-profile of $2\theta_{\infty}$ for M87*.
(a) Refer to caption (b)
Figure 8: The $A_{0}$-profile of the theoretical shadow angular diameter ($2\theta_{\infty}$) of the LGT black hole, within the observed shadow diameters of (a) M87* and (b) Sgr A*, at the $1\sigma$ confidence level.
For Sgr A*, we consider the averaged shadow angular diameter, derived from independent algorithms analyzing EHT observations. These studies suggest that $\theta_{d}$ for Sgr A* lies within the range $46.9\,\mu\text{as}$ to $50\,\mu\text{as}$. Including the $1\sigma$ uncertainty, this range extends from $41.7\,\mu\text{as}$ to $55.6\,\mu\text{as}$ [The Event Horizon Telescope Collaboration] et al. (2022b). In Fig. 8(b), we incorporate this confidence interval to compare the theoretical shadow diameter with the observational results, thereby constraining $A_{0}$. Our findings suggest that the LGT black hole aligns most closely with the EHT data when $A_{0}$ approaches the Schwarzschild limit, as $A_{0}$ remains nearly unity within its uncertainty range.
VII Conclusion
In this work, we have investigated black hole solutions and their observational signatures within the framework of Lorentz gauge gravity, an intriguing alternative to General Relativity motivated by recent cosmological observations and foundational issues in quantum gravity. We have systematically studied the gravitational lensing effects and shadow structures of static, spherically symmetric black hole solutions arising from this theory.
Through explicit analytical and numerical analyses, we have shown that the Lorentz gauge parameter $A_{0}$, characterizing the scale of the connection, significantly influences black hole properties, including horizon radius, photon sphere radius, shadow radius, and gravitational lensing angles. Notably, deviations from the Schwarzschild geometry emerge clearly through variations in $A_{0}$, potentially providing observational discriminators to test the validity of this gravitational framework.
Specifically, we computed weak deflection angles using both Gauss-Bonnet theorem and geodesic methods, confirming consistency between approaches and highlighting modifications introduced by the LGT. In the strong lensing regime, we derived the critical impact parameter and characterized the logarithmic divergence of the deflection angle, identifying clear signatures distinguishing these solutions from classical Schwarzschild black holes.
Furthermore, our analysis of the black hole shadow revealed its explicit dependence on $A_{0}$, exhibiting a pronounced sensitivity that could be tested with high-precision observations such as those from the Event Horizon Telescope. The computed shadow radius provides robust criteria for distinguishing Lorentz gauge gravity from standard General Relativity through direct imaging. These findings demonstrate the potential of gravitational lensing and shadow imaging as powerful observational tools to probe modified gravity theories. Our results thus contribute significantly toward establishing precise observational tests that can discriminate between LGT and classical GR, paving the way for deeper insights into the fundamental nature of gravity and spacetime structure.
Acknowledgements
A.Ö. would like to acknowledge the contribution of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG) and the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology). We also thank EMU, TUBITAK, ULAKBIM (Turkiye) and SCOAP3 (Switzerland) for their support. The work of M.F. is supported by Universidad Central de Chile through the project No. PDUCEN20240008.